In his neighborhood, a businessman would be setting up a repair shop. At a cash price of P 2.8 Million, he would be getting a complete set of tools and equipment from a vendor. If by installment, he would need to pay today P 280,000 and the balance is payable at the end of each quarter for 5 years. The interest is at 18% cpd. qtr. Determine the following: a) the quarterly payment; b) If he chooses to pay at the end of each year for 5 years, how much should it pay?; and c) The value of the investment on the repair shop at the end of 5 years.
Answer #1

(a)

Number of quarters = 4 x 5 = 20

Quarterly (nominal) interest rate = 18%/4 = 4.5%

Quarterly payment ($) = 2,800,000 / P/A(4.5%, 20) = 2,800,000 / 13.0079** = 215,253.81

**P/A(4.5%, 20) = [1 - (1.045)-20] / 0.045 = (1 - 0.4146) / 0.045 = 0.5853 / 0.045 = 13.0079

(b)

Annual payment ($) = 2,800,000 / P/A(18%, 5) = 2,800,000 / 3.1272** = 895,369.66

**From P/A factor table

(b)

(i) Investment value after 5 years for quarterly payments ($) = 215,253.81 x 20 = 4,305,076.2

(ii) Investment value after 5 years for monthly payments ($) = 895,369.66 x 5 = 4,476,848.3